OpenStudy (anonymous):

A robust first grader goes for swimming lessons three times a week, and he enjoys it thoroughly. During the first week of December, he develops an upper respiratory infection, starts coughing continuously, and has a high fever of 39°C. His condition remains the same even after five days of bed rest and conservative measures. Eventually, his mother decides to take him to the doctor who diagnoses it as an acute case of middle ear infection (otitis media), and prescribes amoxicillin for his treatment. He weighs 44lb, and the daily dose of amoxicillin for the treatment of the ear infecti

OpenStudy (anonymous):

1) Interconversion of Celsius to Fahrenheit (F-32) / 9 = C/ 5 Approximation, (C x 2) + 32 = F Thus, temperature in Fahrenheit = 100 2) 1 kg = 2.2 lbs 44 lbs = 20 kg Dose of Amoxicillin = 20 kg x 30 mg / kg = 600 mg per day or 300 mg per dose (i.e. 12 hourly)

OpenStudy (unklerhaukus):

37°C = 102.2°F 44lbs = 19.95kg looks good to me